# Lie symmetry analysis of the two-Higgs-doublet model field equations

### v1

@@
We apply Lie symmetry analysis of partial differential equations (PDEs) to the Euler-Lagrange equations of the two-Higgs-doublet model (2HDM), to determine its scalar Lie point symmetries. A Lie point symmetry is a structure-preserving transformation of the spacetime variables and the fields of the model, which is also continuous and connected to the identity. Symmetries of PDEs may in general be divided into strict variational symmetries, divergence symmetries and non-variational symmetries, where the first two are collectively referred to as variational symmetries. Variational symmetries are usually preserved under quantization, and variational Lie symmetries yield conservation laws. We demonstrate that there are no scalar Lie point divergence symmetries or non-variational Lie point symmetries in the 2HDM, and re-derive its well-known strict variational Lie point symmetries, thus confirming the consistency of our implementation of Lie’s method. Moreover, we prove three general results which may simplify Lie symmetry calculations for a wide class of particle physics models. Lie symmetry analysis of PDEs is a broadly applicable method for determining Lie symmetries. As demonstrated here by example, it can be applied to models with many variables, parameters and reparametrization freedom, while any missing discrete symmetries may be identified through the automorphism groups of the resulting Lie symmetry algebras.
@@ 1 Introduction
Theories with extended scalar sectors beyond the Standard Model (SM) Higgs boson, discovered at CERN in 2012 [1, 2], may explain the matter-antimatter asymmetry of the Universe [3, 4, 5] and accommodate dark matter candidates [6, 7, 8, 9]. New scalars fields can be introduced in several ways, but the most phenomenologically viable and studied implementations are additional complex, scalar $\mathsf{SU}(2)_{L}$ doublets or scalar gauge singlets [10, 11, 12]. The introduction of such fields greatly enlarges the number of viable particle physics models. One of the most important properties of a model, is its symmetry: A symmetry is a structure-preserving transformation of objects such as equations, Lagrangians and action integrals. Symmetries may imply conservation laws, relate apparently distinct models, reduce the number of free parameters, protect parameters from large quantum corrections, ensure parameter relations are stable under quantum corrections and stabilize dark matter candidates against decay.
For multi-Higgs-doublet models (NHDMs), symmetry properties have been studied extensively. The simplest of such models, the two-Higgs-doublet model (2HDM) has received particular attention. The model may facilitate $\operatorname{CP}$ violation [13], necessary for generating the observed matter-antimatter asymmetry during baryogenesis [3], dark matter [14, 15, 7] and is also a necessary part of (minimal) supersymmetric theories [16, 17, 18].
@@
Continuous symmetries connected to the identity, also known as Lie symmetries, may be studied in general by a method originally introduced by Norwegian mathematician Sophus Lie [19], and subsequently developed by his successors. The method, Lie symmetry analysis of partial differential equations (PDEs), can be utilized to find all infinitesimal (sometimes called ”local”) symmetries of a system of PDEs $\Delta$, which correspond to the Lie symmetry algebra $\mathfrak{g}$ of $\Delta$. This method is the original context were the concepts now known as Lie algebras and Lie groups emerged [19]. The Lie symmetry algebra $\mathfrak{g}$ of $\Delta$ determines to a large extent the maximal global symmetry group $\mathsf{G}$ of $\Delta$, since the Lie algebra of $\mathsf{G}$ must equal $\mathfrak{g}$. Lie symmetry analysis may be applied on almost any kind of system of PDEs, e.g. a system of Euler-Lagrange equations $E(\mathcal{L})=0$ for a theory given by a Lagrangian $\mathcal{L}$. When the dependent variables of $\mathcal{L}$ are fields, the Euler-Lagrange equations are also referred to as the field equations of the field theory given by $\mathcal{L}$. Particle physics models like multi-Higgs models are examples of field theories given by Lagrangians.
@@
for each field (i.e. dependent variable) $y^{i}$ and derivative $y^{i}_{,\mu}\equiv\partial y^{i}/\partial x^{\mu}$ occurring in $\mathcal{L}$. Then there are three types of symmetries of the field equations (1.2): First, the strict variational symmetries (SVSs) which leave the action $\mathcal{S}$ strictly invariant, i.e.
@@
is the integral of a total divergence $d_{\mu}\beta^{\mu}$, where $d_{\mu}\equiv d/dx^{\mu}$ is a total derivative, for non-vanishing, local functions $\beta^{\mu}$. Here $\Delta\mathcal{S}$ is, by the divergence theorem, converted to a boundary term, which vanishes under the usual boundary conditions on the fields, as shown in the last equality of (1.4). These first two types of symmetries are the variational symmetries, symmetries of the action $\mathcal{S}$. Variational symmetries usually translate to symmetries of the corresponding quantum theory, provided the measure of Feynman’s path integral is invariant as well as the action $\mathcal{S}$. Moreover, variational Lie symmetries lead to conserved currents by Noether’s theorem [20]. The third kind of symmetries are the non-variational symmetries that preserve the structure of the field equations, but do not leave the action $\mathcal{S}$ invariant.
@@
When a symmetry only transforms free and dependent variables $(x,y)$, it is called a point symmetry. The derivatives of the dependent variables are also transformed under a point symmetry, but these transformations are dictated by the transformations of the variables $(x,y)$. This means a point symmetry cannot interchange a dependent variable with its derivative.
@@
The possible, inequivalent (real) scalar point SVS groups of the 2HDM was determined in [21], using a formalism of gauge invariant, scalar bilinears. A new, discrete point transformation in the 2HDM which leads to parameter relations respected by quantum corrections, characteristic for variational symmetries, was identified in [22]. As suggested in [22], this transformation may be regarded as a complex, discrete symmetry. Such transformations may be a generic feature of quantum field theories, although they do not necessarily have to be interpreted as complex symmetries [23]. However, complex symmetries of real structures are not unknown in the context of symmetry analysis of differential equations, see e.g. [24] and [25] for examples of complex, discrete symmetries of real differential equations. An overview of the 2HDM may be found in [26]. The possible scalar, point SVSs of the 3HDM were classified in [27, 28] while the categorization of scalar symmetry groups in the 4HDM has recently been initiated in [29, 30]. In these studies, finite group theory was applied as a main tool. The possible scalar, point SVSs of the 2HDM and 3HDM, including custodial symmetries, were determined in [31, 32] and [33], respectively, by applying bilinear and tensor product methods. The scalar, Lie point SVSs of the general NHDM’s kinetic terms were determined in [34], including the corresponding symmetries in the custodial limit.
@@
The purpose of this work is twofold: First, inspired by the aforementioned discovery in [22] of a new, complex 2HDM symmetry, we want to investigate the possibility of having Lie point divergence and non-variational symmetries in a 2HDM. Since all Higgs doublets carry the same quantum numbers (that is, the same isospin and hypercharge), unitary linear combinations of the original doublets, accompanied by appropriate redefinitions of parameters to keep the Lagrangian invariant, relate different parametrizations that describe the same physics. This phenomenon is known as basis freedom or reparametrization freedom. We will in our analysis choose specific bases were certain parameters are vanishing, to, as much as practically possible, prevent the same symmetry from occurring in multiple bases. Second, we would like to demonstrate how to classify all Lie point symmetries in models with many variables, parameters and reparametrization freedom, like the 2HDM. Both of these aims will be pursued by employing Lie symmetry analysis to the field equations of the general 2HDM.
@@
Sections 2.1-2.3 review the relevant theory of Lie symmetry analysis of PDEs, while we present three new results in Sections 2.4 and 2.5, where the first result, Theorem 1, will be crucial for Section 4.3. We then give a simple, concrete example of how the theory of Sections 2.1-2.5 may be applied in Section 3. Section 3 is not necessary for reading Section 4, and may be skipped by readers familiar with Lie symmetry analysis of PDEs or readers who want to go straight to the main application in Section 4. In Section 4 we recall standard formalism of the 2HDM, and in Sections 4.2-4.3 we perform the Lie symmetry analysis of the 2HDM. A summary and outlook is given in Section 5, while a proof of a result of Section 2.5, Proposition 2.5, is delegated to Appendix A.
@@
In this article we adopt the mathematicians’ convention for Lie algebras, which means that a matrix Lie group $\mathsf{G}$ with Lie algebra $\mathfrak{g}$ is generated by $\exp(\mathfrak{g})$, and not by $\exp(i\mathfrak{g})$ which would correspond to the physicists’ convention. Moreover, $d_{\mu}\equiv d/dx^{\mu}$ denotes the total derivative, while $D_{\mu}$ is reserved for the covariant derivative (2.52), unless the index is a multi-index $J$, see section 2.2, or a function $P$, implying a Fréchet derivative $D_{P}$. Repeated indices are implicitly summed over, while a check over the indices imply the summation convention is dispensed so that, for instance, $\eta^{\widecheck{\imath}}\partial_{\phi_{\widecheck{\imath}}}$ only consists of one term. Finally, we will use respectively $\Re$ and $\Im$ for real and imaginary parts of expressions, e.g. will $\lambda_{5}=\Re(\lambda_{5})+i\,\Im(\lambda_{5})$.
@@ 2 Lie symmetry analysis of PDEs
In this section we will give an overview of relevant Lie symmetry theory of PDEs. Most of the material in Sections 2.1-2.3 can be found in [35, 36, 24, 37, 38].
@@ 2.1 Point symmetries of systems of PDEs
Consider a $n$’th order system of PDEs given by
@@
with $d$ independent variables $x=(x^{0},\ldots,x^{d-1})$ (spacetime coordinates) and $q$ dependent variables $y=(y^{1},\ldots,y^{q})$ (the fields), where $y^{(k)}$ denotes derivatives of order $k$ of the dependent variables $y^{j}$ with respect to the independent variables $x^{\mu}$. We abbreviate (2.1) by $\Delta=0$ or just $\Delta$.
Then, a point symmetry $S$ of the system of PDEs (2.1) is a diffeomorphism [^1] Which means $S$ and $S^{-1}$ exist and have at least continuous first order derivatives, but are often taken to be smooth, i.e $C^{\infty}$. on the space of variables (i.e. it is a point transformation), which maps solutions of (2.1) to solutions. More precisely, this means
@@
holds whenever (2.1) holds. In (2.3) the action of $S$ is prolonged to the derivatives such that they are mapped to the corresponding derivatives in the transformed variables, that is
@@ 2.2 Infinitesimal generators and their prolongations
for an infinitesimal $\epsilon$. The $k$-prolongation of $X$, $\operatorname{pr}^{(k)}X$, extends this action to derivatives up to $k$’th order,
@@
by a formal, infinite series (2.8) with $k=\infty$. If $\xi^{\mu}=0$ for all $\mu$ the infinitesimal point transformation generator $X$ is said to be in evolutionary form, and for a general $X$ the infinitesimal generator
@@
is called its evolutionary representative. In case $Q$ includes derivatives, $X_{Q}$ will be a so-called generalized symmetry of the system of PDEs if $X$ is a symmetry thereof [35], but we will only consider cases where $X_{Q}$ is a point symmetry due to $\xi=0$ in (2.13). For a point transformation given by $X$ already in evolutionary form ($\xi=0$) its prolongation simply becomes
@@
In case of a very simple generator $X=y\partial_{y}$ and only one independent variable $x$ the 1-prolongation $\operatorname{pr}^{(1)}(X)=y\partial_{y}+y^{\prime}\partial_{y^{\prime}}$, and it hence just extends the infinitesimal transformation $y\to(1+\epsilon X)y=y+\epsilon y$ to the first order derivative, since $y^{\prime}\to(1+\epsilon\operatorname{pr}^{(1)}(X))y^{\prime}=y^{\prime}+\epsilon y^{\prime}$. If independent variables also transform, i.e. $\xi^{\mu}\neq 0$ for some $\mu$, the prolongations may become much more complicated, as testified by the formulas above.
Now, let $\mathsf{G}$ be a connected Lie group and let $\mathfrak{g}$ be its Lie algebra. Then $\mathsf{G}$ is a symmetry group [^2] Possibly a local group, i.e. only defined close to identity. of a fully regular333”Fully regular” here means the system $\Delta$ is locally solvable and have a non-vanishing Jacobi determinant. In practice, most systems are fully regular [36]. system of $m$ PDEs given by $\Delta=0$ if and only if
@@
[…]
The condition (2.17) is known as the linearized symmetry condition and the resulting, concrete equations are called the determining equations of $\mathfrak{g}$ (or, less precise, $\mathsf{G}$). Combining (2.8) and (2.17) shows that the set of determining equations yields an extensive, over-determined system of linear PDEs in the coefficients $\xi^{\mu}(x,y)$ and $\eta^{i}(x,y)$. This system may almost always be explicitly solved, and hence we can determine the symmetry algebra $\mathfrak{g}$, i.e. the infinitesimal symmetries. In this work we, most of the time, will apply the Mathematica package SYM [45] to calculate determining equations.

### v2

@@
We apply Lie symmetry analysis of partial differential equations (PDEs) to the Euler-Lagrange equations of the two-Higgs-doublet model (2HDM), to determine its scalar Lie point symmetries. A Lie point symmetry is a structure-preserving transformation of the spacetime variables and the fields of the model, which is also continuous and connected to the identity. Symmetries of PDEs may, in general, be divided into strict variational, divergence and non-variational symmetries, where the first two are collectively referred to as variational symmetries. Variational symmetries are usually preserved under quantization, and variational Lie symmetries yield conservation laws. We demonstrate that there are no scalar Lie point divergence symmetries or non-variational Lie point symmetries in the 2HDM, and re-derive its well-known strict variational Lie point symmetries, thus confirming the consistency of our implementation of Lie’s method. Moreover, we prove three general results that may simplify Lie symmetry calculations for a wide class of particle physics models. Lie symmetry analysis of PDEs is broadly applicable for determining Lie symmetries. As demonstrated in this work, the method can be applied to models with many variables, parameters, and reparametrization freedom, while any missing discrete symmetries can be identified through the automorphism groups of the resulting Lie symmetry algebras.
@@ 1 Introduction
The discovery of the Standard Model (SM) Higgs boson at the LHC in 2012 by the ATLAS [1] and CMS [2] collaborations has motivated a broad exploration of theories with extended scalar sectors. Such extensions may help explain the observed matter-antimatter asymmetry of the universe, starting from Sakharov’s conditions for baryogenesis [3] and their possible realization in electroweak baryogenesis scenarios. In particular, multi-Higgs-doublet models such as two-Higgs-doublet models (2HDMs) [4] and three-Higgs-doublet models (3HDMs) [5] can provide additional sources of $\operatorname{CP}$ violation and may strengthen the electroweak phase transition.
Extended scalar sectors can also provide viable dark matter candidates. Gauge singlet scalars offer a minimal realization [6], whereas inert-doublet models provide an alternative framework [7]. The inert-doublet scenario has been analysed in detail in [8], while associated gamma-ray line signatures were investigated in [9]. Additional realizations arise in mixed doublet–singlet models, such as singlet-extended 2HDMs [10], and in multi-doublet constructions with stabilizing symmetries, e.g. $\mathsf{U}(1)$-stabilized 3HDMs [11]. As indicated, phenomenological studies have largely focused on additional complex $\mathsf{SU}(2)_{L}$ doublets [12] and on scalar gauge singlets. A comprehensive textbook treatment is given in [13], while a more recent review can be found in [14].
@@
The introduction of such fields greatly enlarges the number of viable particle physics models. One of the most important characteristics of a model is its set of symmetries. By a symmetry we mean a structure-preserving transformation of objects such as field equations, Lagrangians, and action integrals. Symmetries may imply conservation laws, relate apparently distinct models, reduce the number of free parameters, protect parameters from large quantum corrections, ensure that parameter relations are stable under quantum corrections and stabilize dark matter candidates against decay.
Symmetry properties have been studied extensively for multi-Higgs-doublet models (NHDMs), with the 2HDM as the simplest and most widely analysed example. In particular, the 2HDM can accommodate $\operatorname{CP}$ violation in the scalar sector [15], making it a natural framework for scenarios of electroweak baryogenesis. It also plays a central role in dark matter model building and in supersymmetric extensions of the SM. A two-Higgs-doublet structure is, for example, an essential ingredient of minimal supersymmetric models; see [16] and the subsequent discussion in [17], as well as the detailed analysis in [18].
@@
Continuous symmetries connected to the identity, also known as Lie symmetries, may be studied in general by a method originally introduced by Norwegian mathematician Sophus Lie [19], and subsequently developed by his successors. The method, Lie symmetry analysis of partial differential equations (PDEs), can be utilized to find all infinitesimal (sometimes called ”local”) symmetries of a system of PDEs $\Delta$, which correspond to the Lie symmetry algebra $\mathfrak{g}$ of $\Delta$. This method is the original context in which the concepts now known as Lie algebras and Lie groups emerged [19]. The Lie symmetry algebra $\mathfrak{g}$ of $\Delta$ determines, to a large extent, the maximal global symmetry group $\mathsf{G}$ of $\Delta$, because the Lie algebra of $\mathsf{G}$ must equal $\mathfrak{g}$. Lie symmetry analysis may be applied to almost any kind of system of PDEs, for example a system of Euler-Lagrange equations $E(\mathcal{L})=0$ for a theory given by a Lagrangian $\mathcal{L}$. When the dependent variables of $\mathcal{L}$ are fields, the Euler-Lagrange equations are also referred to as the field equations of the field theory given by $\mathcal{L}$. Particle physics models, such as multi-Higgs models, are examples of field theories given by Lagrangians.
@@
for each field (i.e. dependent variable) $y^{i}$ and derivative $y^{i}_{,\mu}\equiv\partial y^{i}/\partial x^{\mu}$ occurring in $\mathcal{L}$, where $d_{\mu}\equiv d/dx^{\mu}$ is a total derivative. Then, there are three types of symmetries of the field equations (1.2): First, the strict variational symmetries (SVSs) that leave the action $\mathcal{S}$ strictly invariant, that is,
@@
is the integral of a total divergence $d_{\mu}\beta^{\mu}$, with non-vanishing, local functions $\beta^{\mu}$. Here, $\Delta\mathcal{S}$ is, by the divergence theorem, converted to a boundary term, which vanishes under the usual boundary conditions on the fields, as shown in the last equality of (1.4). These first two types of symmetries are the variational symmetries, symmetries of the action $\mathcal{S}$. Variational symmetries are typically preserved in the corresponding quantum theory, provided both the action $\mathcal{S}$ and the measure of Feynman’s path integral remain invariant. Moreover, variational Lie symmetries lead to conserved currents by Noether’s theorem [20]. The third type of symmetries are the non-variational symmetries that preserve the structure of the field equations, but do not leave the action $\mathcal{S}$ invariant.
@@
When a symmetry only transforms free and dependent variables $(x,y)$, it is called a point symmetry. The derivatives of the dependent variables are also transformed under a point symmetry, however, these transformations are dictated by the transformations of the variables $(x,y)$. This implies that a point symmetry cannot, for instance, interchange a dependent variable with its derivative.
@@
The possible, inequivalent (real) scalar point SVS groups of the 2HDM were determined in [21], using a formalism of gauge invariant scalar bilinears. A new discrete point transformation in the 2HDM that leads to parameter relations respected by quantum corrections, characteristic of variational symmetries, was identified in [22]. As suggested in [22], this transformation may be regarded as a complex, discrete symmetry. Such transformations may be a generic feature of quantum field theories, although they do not necessarily need to be interpreted as complex symmetries [23]. However, complex symmetries of real structures are not unknown in the context of symmetry analysis of differential equations, see e.g. [24] or [25] for examples of complex, discrete symmetries of real differential equations. An overview of the 2HDM can be found in [26]. The possible scalar point SVSs of the 3HDM have been classified in stages: Abelian symmetry groups were analysed in [27], and the full classification was completed in [28]. The corresponding categorization of scalar symmetry groups in the 4HDM has only recently been initiated, with extensions of cyclic groups studied in [29] and extensions by rephasing groups in [30]. In these works, finite group theory plays a central role.
The scalar point SVSs of the 2HDM and 3HDM, including custodial symmetries, have also been classified. For the 2HDM, this programme was initiated in [31] and completed in [32]. For the 3HDM, the corresponding classification was obtained in [33], using bilinear and tensor-product methods. Finally, the scalar, Lie point SVSs of the general NHDM kinetic terms were determined in [34], including the corresponding symmetries in the custodial limit.
@@
The purpose of this work is twofold: First, inspired by the aforementioned discovery in [22] of a new, complex 2HDM symmetry, we want to investigate the possibility of having Lie point divergence and non-variational symmetries in a 2HDM. Because all Higgs doublets carry the same quantum numbers (i.e., the same isospin and hypercharge), unitary linear combinations of the original doublets, accompanied by appropriate redefinitions of parameters to keep the Lagrangian invariant, relate different parameterizations that describe the same physics. This phenomenon is known as basis freedom or reparametrization freedom. In our analysis, we will choose specific bases where certain parameters vanish, in order to minimize the occurrence of equivalent symmetries across different bases. Second, we would like to demonstrate how to classify all Lie point symmetries in models with many variables, parameters and reparametrization freedom, like the 2HDM. Both of these aims will be pursued by applying Lie symmetry analysis to the field equations of the general 2HDM.
@@
Sections 2.1-2.3 review the relevant theory of Lie symmetry analysis of PDEs, while we present three new results in Sections 2.4 and 2.5, where the first result, Theorem 1, is crucial for Section 4.3. We then provide a simple, concrete example of how the theory of Sections 2.1-2.5 may be applied in Section 3. Section 3 is not necessary for reading Section 4, and may be skipped by readers familiar with Lie symmetry analysis of PDEs or readers who want to go straight to the main application in Section 4. In Section 4, we recall standard formalism of the 2HDM, and in Sections 4.2-4.3 we perform the Lie symmetry analysis of the 2HDM. A summary and outlook is given in Section 5, while a proof of a result of Section 2.5, Proposition 2.5, is delegated to Appendix A.
@@
In this article we adopt the mathematicians’ convention for Lie algebras, which means that a matrix Lie group $\mathsf{G}$ with Lie algebra $\mathfrak{g}$ is generated by $\exp(\mathfrak{g})$, and not by $\exp(i\mathfrak{g})$ which would correspond to the physicists’ convention. Moreover, $d_{\mu}\equiv d/dx^{\mu}$ denotes the total derivative, whereas $D_{\mu}$ is reserved for the covariant derivative (2.55), unless the index is a multi-index $J$, see Section 2.2, or a function $P$, implying a Fréchet derivative $D_{P}$. Repeated indices are implicitly summed over, while a check over the indices implies that the summation convention is dispensed so that, for instance, $\eta^{\widecheck{\imath}}\partial_{\phi_{\widecheck{\imath}}}$ only consists of one term. Finally, we will use respectively $\Re$ and $\Im$ for real and imaginary parts of expressions, e.g. will $\lambda_{5}=\Re(\lambda_{5})+i\,\Im(\lambda_{5})$.
@@ 2 Lie symmetry analysis of PDEs
In this section, we provide a brief overview of the Lie symmetry theory for systems of PDEs. A classical introduction is given by Olver [35], with a more concise account in his lecture notes [36]. Further treatments are available in the textbooks by Hydon [24], Bluman and Kumei [37], and Cantwell [38].
@@ 2.1 Point symmetries of systems of PDEs
Consider an $n$th-order system of PDEs given by
@@
with $d$ independent variables $x=(x^{0},\ldots,x^{d-1})$ (spacetime coordinates) and $q$ dependent variables $y=(y^{1},\ldots,y^{q})$ (the fields), where $y^{(k)}$ denotes derivatives of order $k$ of the dependent variables $y^{j}$ with respect to the independent variables $x^{\mu}$. We abbreviate (2.1) as $\Delta=0$ or just $\Delta$.
Then, a point symmetry $S$ of the system of PDEs (2.1) is a diffeomorphism [^1] Which means that $S$ and $S^{-1}$ exist and have at least continuous first-order derivatives, but are often taken to be smooth, that is, $C^{\infty}$. on the space of variables (i.e., it is a point transformation), which maps solutions of (2.1) to solutions. More precisely, this means
@@
holds whenever equation (2.1) holds. In (2.3), the action of $S$ is prolonged to the derivatives, such that they are mapped to the corresponding derivatives in the transformed variables, that is,
@@
The expressions $\Delta_{i}$ in (2.1) can also be regarded as functions on the $n$th jet space $J^{n}$, which in the present setting can be identified with Euclidean space $\mathbb{R}^{s}$ with formal coordinates $(x,y,\ldots,y^{(n)})$ corresponding to all independent and dependent variables and their distinct derivatives up to order $n$. This implies $s=d+q\binom{d+n}{n}$. Each PDE can then be viewed as a smooth map
$\Delta_{i}:J^{n}\to\mathbb{R},$ (2.6)
and the system (2.1) can be described by the solution submanifold $\mathcal{M}_{\Delta}\subset J^{n}$,
$\mathcal{M}_{\Delta}=\bigl\{(x,y,\ldots,y^{(n)})\in J^{n}\;\big|\;\Delta_{i}(x,y,\ldots,y^{(n)})=0\ \text{for all }i\bigr\},$ (2.7)
which consists of all points $(x,y,\ldots,y^{(n)})\in J^{n}$ that satisfy the system.
@@ 2.2 Infinitesimal generators and their prolongations
for an infinitesimal $\epsilon$. The $k$th prolongation of $X$, denoted $\operatorname{pr}^{(k)}X$, is the vector field on the $k$th jet space $J^{k}$ obtained by extending the action of $X$ to the derivatives of the dependent variables, namely
@@
as the formal vector field obtained by extending (2.10) to all derivatives of arbitrary order, so that $\operatorname{pr}X$ is the extension of $X$ to the infinite jet space $J^{\infty}$. If $\xi^{\mu}=0$ for all $\mu$ the infinitesimal point transformation generator $X$ is said to be in evolutionary form, and for a general $X$ the infinitesimal generator
@@
is called its evolutionary representative. In the case $Q$ includes derivatives, $X_{Q}$ will be a so-called generalized symmetry of the system of PDEs if $X$ is a symmetry thereof [35], but we will only consider cases where $X_{Q}$ is a point symmetry due to $\xi=0$ in (2.15). For a point transformation given by $X$ already in evolutionary form ($\xi=0$), its prolongation simply becomes
@@
In case of a very simple generator $X=y\partial_{y}$ and only one independent variable $x$ the 1-prolongation $\operatorname{pr}^{(1)}(X)=y\partial_{y}+y^{\prime}\partial_{y^{\prime}}$, and it hence just extends the infinitesimal transformation $y\to(1+\epsilon X)y=y+\epsilon y$ to the first-order derivative, since $y^{\prime}\to(1+\epsilon\operatorname{pr}^{(1)}(X))y^{\prime}=y^{\prime}+\epsilon y^{\prime}$. If the independent variables also transform, that is, $\xi^{\mu}\neq 0$ for some $\mu$, the prolongations may become much more complicated, as testified by the formulas above.
Now, let $\mathsf{G}$ be a connected Lie group acting (locally) on $\mathbb{R}^{d+q}$ with coordinates $(x,y)$, where $x\in\mathbb{R}^{d}$ are the independent variables and $y\in\mathbb{R}^{q}$ the dependent variables, and let $\mathfrak{g}$ be its Lie algebra. Then $\mathsf{G}$ is a symmetry group [^2] Possibly a local group, i.e. only defined in a neighbourhood of the identity. of a fully regular333“Fully regular” here means that the system $\Delta$ is locally solvable and has a non-vanishing Jacobian determinant. In practice, most systems are fully regular [36]. system of $m$ PDEs, written as $\Delta=0$, if and only if
@@
for all infinitesimal generators $X\in\mathfrak{g}$ [36]. The condition (2.19) is equivalent to requiring that, for a system $\Delta$ of order $n$, the prolongation $\operatorname{pr}^{(n)}X$ (or $\operatorname{pr}X$) is a vector field tangent to the solution manifold $\mathcal{M}_{\Delta}$ defined in (2.7), which can also be taken as a geometric definition of a (Lie point) symmetry.
The elements of $\mathsf{G}$ are (Lie point) symmetries of the system $\Delta$. When no confusion can arise, we will also refer to a generator $X\in\mathfrak{g}$ as a symmetry whenever it satisfies (2.19). Moreover, all Lie point symmetries of a system of PDEs will be found by Lie symmetry analysis, because (2.19) holds for all symmetry generators in any connected symmetry group. We will only consider point symmetries, implemented by only applying point transformation generators (2.8) in (2.19), but (2.19) may be generalized to higher-order symmetries [36]. Of course, we only have to consider prolongations $\operatorname{pr}^{(n)}(X)$ up to the order $n$ of $\Delta$, when applying (2.19). If we perform a Lie symmetry analysis of the system $\Delta=0$ we may for example find that the Lie symmetry algebra is $\mathfrak{g}=\mathfrak{so}(6)$. Then, a (global) symmetry group of the system $\Delta=0$ may be $\mathsf{SO}(6)$. However, if I-I acts identically to $I$ on $\Delta$, the corresponding faithful symmetry group is the projective group $\mathsf{PSO}(6)\cong\mathsf{SO}(6)/\{\pm I\}$. Both groups have the same Lie algebra $\mathfrak{g}$. However, the maximal, faithfully (or effectively) acting symmetry group may be a larger group than the latter, such as $\mathsf{PO}(6)=\mathsf{O}(6)/\{\pm I\}$, also with Lie algebra $\mathfrak{g}$, but with two components, due to the presence of an additional discrete reflection symmetry. Generally, if the symmetry algebra is $\mathfrak{g}$, the maximal symmetry group of $\Delta$ may, a priori, be any group $\mathsf{G}$ with Lie algebra $\mathfrak{g}$. The identity component of $\mathsf{G}$ is then a quotient $\tilde{\mathsf{G}}/N$ of the unique simply connected group $\tilde{\mathsf{G}}$ with Lie algebra $\mathfrak{g}$, where $N$ is a normal subgroup [39]. The identification of the maximal symmetry group $\mathsf{G}$, including any missing discrete symmetries, can be achieved by studying the automorphism groups of the Lie symmetry algebras obtained through Lie symmetry analysis. For ordinary differential equations, Hydon [40] describes how discrete point symmetries can be constructed from the continuous symmetry algebra, and extends this approach to discrete contact symmetries in [41], while in [42] similar methods are applied to partial differential equations. A systematic application of these techniques to determine the full symmetry group $\mathsf{G}$ for the models considered here would, however, require a separate analysis and is beyond the scope of the present work.
Condition (2.19) is known as the linearized symmetry condition, and the resulting concrete equations are called the determining equations of $\mathfrak{g}$ (or, less precisely, $\mathsf{G}$). Combining equations (2.10) and (2.19) shows that the set of determining equations yields an extensive, over-determined system of linear PDEs in the coefficients $\xi^{\mu}(x,y)$ and $\eta^{i}(x,y)$. This system may almost always be explicitly solved, and hence we can determine the symmetry algebra $\mathfrak{g}$, i.e. the infinitesimal symmetries. In this work we, most of the time, will apply the Mathematica package SYM [43] to calculate determining equations.